# Non-Gaussian state preparation and enhancement using weak-value amplification

### v1

@@ IV Enlargement of Schrödinger cat states
The amplitude-enlarging advantage of our protocol for SC states can also be confirmed by using the even SC state as an example. Similar to the odd SC case, our protocol can generate large-amplitude even SC states with high fidelity when a small-amplitude even SC state is used as input. As shown in Fig. 10, when the input even SC states have amplitudes $\alpha=1.5$, $2.0$, and $3.0$, the corresponding enlarged even SC states with amplitudes $\alpha^{\prime}=2.0$, $2.41$, and $3.3$ are obtained, with fidelities $F_{+}=0.950$, $0.981$, and $0.995$, respectively. These enlarged states are encoded in the output state $|\Psi_{+}\rangle$, generated by our protocol for a given ideal small-amplitude input, even SC state. As shown in Fig. 8(d), the amplitude enlargement ratio $\eta_{+}=\alpha^{\prime}/\alpha$ for even SC states remains greater than one for all tested values of $\alpha$, further confirming that our protocol exhibits a robust amplitude enhancing effect for even SC states. Moreover, the even SC state closely resembles a Gaussian squeezed vacuum state when $\alpha\lessapprox 0.75$, and in that regime, it serves as a high-fidelity ($F=0.99$) approximation to the SV state (Brańczyk and Ralph, 2008). Therefore, for even SC inputs with $\alpha\lessapprox 0.75$, the output behavior mirrors the results discussed in Sec. III.2, using the optimized squeezing parameter $r_{opt}(\alpha)=\ln\left(\sqrt{2\alpha^{2}+\sqrt{1+4\alpha^{2}}}\right)$. Although the even SC state is approximately Gaussian in this small-amplitude regime, the corresponding output state generated by our protocol displays strong nG features, as indicated by the appearance of Wigner negativity [see Figs. 5(c) and 5(d)]. The amplitude enlargement effect of our protocol can also be directly confirmed by comparing the average photon numbers of the input and output states. Large-amplitude SC states naturally contain more photons than small-amplitude states. In the output states $|\Psi_{\pm}\rangle$, the first term corresponds to the input SC state, whereas the second term, $b^{\dagger}b|SCS_{\pm}\rangle$, contributes significantly to the photon number. As a result, the average photon number in the output states $|\Psi_{\pm}\rangle$ is always larger than that in the corresponding ideal SC inputs, while still maintaining high fidelity within the discussed parameter regimes.
@@
It should be emphasized that the amplitude-enlarging feature of our protocol for SC states remains valid across a wide range of coherent amplitudes $\alpha$, with the corresponding fidelity $F_{-}$ ($F_{+}$) exceeding 0.99 for $\alpha\gtrsim 2.41$. In these regimes, the enlargement coefficients $\eta_{-}$ ($\eta_{+}$) are also consistently greater than one. The above results indicate that even though we did not directly prepare the exact large-amplitude SC states, the equivalent SC states $|SC_{\pm}(\alpha^{\prime})\rangle$ were encoded in our signal output states $|\Psi_{\pm}\rangle$. Consequently, the two coherent components |±$|\pm\alpha^{\prime}\rangle$ of the equivalent SC states $|SC_{\pm}(\alpha^{\prime})\rangle$ become approximately orthogonal when $\alpha^{\prime}\geq 2$, making such states valuable resources for exploring macroscopic quantum superpositions and quantum information applications.
@@
As discussed in the Introduction, optical SC states can be generated via cross-Kerr interactions characterized by third-order nonlinear susceptibility $\chi^{(3)}$ (Gerry, 1999; Gerry and Knight, 2004). However, in practice, generating large-amplitude SC states under realistic conditions is extremely challenging due to the typically small and attenuated values of $\chi^{(3)}$. As analyzed above, by using the WVA technique to effectively enhance single-photon level nonlinearities, our protocol enables the generation of large-amplitude SC states ($\alpha\geq 2$) with high fidelity, even in the presence of weak cross-Kerr interactions.
@@
In recent years, several methods have been explored for enlarging SC states (Lund et al., 2004; Sychev et al., 2017; Hacker et al., 2019; Li et al., 2024). Among these, the synthesis-based method for enlarging even SC states, as described in Refs. (Lund et al., 2004; Sychev et al., 2017) requires the input to be in a small odd SC state. However, this approach is constrained by the amplitude of the initial state and is resource-intensive, making the generation of large-amplitude SC states experimentally demanding. In Ref. (Hacker et al., 2019), SC states with amplitudes up to $\alpha=1.4$ were generated through direct interaction between a coherent light pulse and a single-sided cavity containing a three-level atom. Although promising, this method is limited in scalability to larger amplitudes, and is susceptible to decoherence when strong light fields are used.
@@
In contrast, our protocol provides a simple and experimentally feasible alternative for generating large-amplitude SC states with high fidelity across a broad range of amplitudes $\alpha$, without requiring explicit photon addition or subtraction operations. However, we note that the success probability of our scheme is not high and is proportional to the square of the beam splitter deviation from the ideal $50:50$ configuration, i.e., $|\epsilon|^{2}$.
V Conclusions
In this study, we propose a feasible protocol for the preparation of the nG quantum state through weak measurement, implementing a weak cross-Kerr interaction characterized by third-order nonlinear susceptibility $\chi^{(3)}$. A single photon was injected into the interference setup through the idler input port. Upon the detection of a single photon at the dark port of the idler beam, the Gaussian input state at the signal port is transformed into a nG state. This transformation is achieved through the amplification of the single-photon nonlinearity, which is enabled by the WVA technique. As specific examples of the protocol, we demonstrate the generation of various nG states, including SPAC states, displaced (squeezed) single- and two-photon states, and TPASV states, all of which have high fidelity.
Furthermore, this protocol enables the generation of a continuous range of nG states with intermediate properties that depend on the initial state of the signal. We also show that the protocol can be employed to enhance the non-Gaussianity of nG states. We find that, using SC states as a test case, the Wigner function negativity, key indicator of non-Gaussianity, significantly increased after passing through our setup. In particular, the protocol also effectively amplifies small SC states into larger ones with coherent amplitudes $\alpha\geq 2$ while maintaining high fidelity ($F>0.99$), particularly in the case of odd SC states.
Throughout the process, the transformation of the Gaussian input into a nG state is primarily owing to an effective sequence of single-photon subtraction followed by addition in the signal mode, facilitated by the weak cross-Kerr interaction and WVA. Remarkably, photon addition and subtraction occur without the need for explicit nG operations. Our results further confirm the utility of postselected weak measurements in quantum state engineering and suggest that the ability to generate a continuous family of nG states may find practical applications in a variety of quantum information processing tasks.
Acknowledgements.

### v2

@@ A Versatile Protocol via Weak Measurement
The amplitude-enlarging advantage of our protocol for SC states can also be confirmed by using the even SC state as an example. Similar to the odd SC case, our protocol can generate large-amplitude even SC states with high fidelity when a small-amplitude even SC state is used as input. As shown in Fig. 10, when the input even SC states have amplitudes $\alpha=1.5$, $2.0$, and $3.0$, the corresponding enlarged even SC states with amplitudes $\alpha^{\prime}=2.0$, $2.41$, and $3.3$ are obtained, with fidelities $F_{+}=0.950$, $0.981$, and $0.995$, respectively. These enlarged states are encoded in the output state $|\Psi_{+}\rangle$, generated by our protocol for a given ideal small-amplitude input even SC state. As shown in Fig. 8d, the amplitude enlargement ratio $\eta_{+}=\alpha^{\prime}/\alpha$ for even SC states remains greater than one for all tested values of $\alpha$, further confirming that our protocol exhibits a robust amplitude enhancing effect for even SC states. Moreover, the even SC state closely resembles a Gaussian squeezed vacuum state when $\alpha\lessapprox 0.75$, and in that regime, it serves as a high-fidelity ($F=0.99$) approximation to the SV state (PhysRevA.78.052304). Therefore, for even SC inputs with $\alpha\lessapprox 0.75$, the output behavior mirrors the earlier analysis for squeezed vacuum inputs, using the optimized squeezing parameter $r_{opt}(\alpha)=\ln\left(\sqrt{2\alpha^{2}+\sqrt{1+4\alpha^{2}}}\right)$. Although the even SC state is approximately Gaussian in this small-amplitude regime, the corresponding output state generated by our protocol displays strong nG features, as indicated by the appearance of Wigner negativity [see Figs. 5c and 5d]. The amplitude enlargement effect of our protocol can also be directly confirmed by comparing the average photon numbers of the input and output states. Large-amplitude SC states naturally contain more photons than small-amplitude states. In the output states $|\Psi_{\pm}\rangle$, the first term corresponds to the input SC state, whereas the second term, $b^{\dagger}b|SCS_{\pm}\rangle$, contributes significantly to the photon number. As a result, the average photon number in the output states $|\Psi_{\pm}\rangle$ is always larger than that in the corresponding ideal SC inputs, while still maintaining high fidelity within the discussed parameter regimes.
@@
It should be emphasized that the amplitude-enlarging feature of our protocol for SC states remains valid across a wide range of coherent amplitudes $\alpha$, with the corresponding fidelity $F_{-}$ ($F_{+}$) exceeding 0.99 for $\alpha\gtrsim 2.41$. In these regimes, the enlargement coefficients $\eta_{-}$ ($\eta_{+}$) are also consistently greater than one. The above results indicate that even though we did not directly prepare the exact large-amplitude SC states, the equivalent SC states $|SC_{\pm}(\alpha^{\prime})\rangle$ were encoded in our signal output states $|\Psi_{\pm}\rangle$. Consequently, the two coherent components $|\pm\alpha^{\prime}\rangle$ of the equivalent SC states $|SC_{\pm}(\alpha^{\prime})\rangle$ become approximately orthogonal when $\alpha^{\prime}\geq 2$, making such states valuable resources for exploring macroscopic quantum superpositions and quantum information applications. However, to maintain the weak nonlinearity condition $g\langle n_{b}\rangle\ll 1$, the coherent amplitudes $\alpha$ of the initial input SC states cannot be too large.
@@
As mentioned in the Introduction, optical SC states can be generated via cross-Kerr interactions characterized by third-order nonlinear susceptibility $\chi^{(3)}$ (PhysRevA.59.4095; Int). However, in practice, generating large-amplitude SC states under realistic conditions is extremely challenging due to the typically small and attenuated values of $\chi^{(3)}$. As analyzed above, by using the WVA technique to effectively enhance single-photon level nonlinearities, our protocol enables the generation of large-amplitude SC states ($\alpha\geq 2$) with high fidelity, even in the presence of weak cross-Kerr interactions.
@@
In recent years, several methods have been explored for enlarging SC states (PhysRevA.70.020101; 2017n; 2019R; 2024Li). Among these, the synthesis-based method for enlarging even SC states, as described in Refs. (PhysRevA.70.020101; 2017n) requires the input to be in a small odd SC state. However, this approach is constrained by the amplitude of the initial state and is resource-intensive, making the generation of large-amplitude SC states experimentally demanding. In Ref. (2019R), SC states with amplitudes up to $\alpha=1.4$ were generated through direct interaction between a coherent light pulse and a single-sided cavity containing a three-level atom. Although promising, this method is limited in scalability to larger amplitudes, and is susceptible to decoherence when strong light fields are used.
@@
In contrast, our protocol provides a simple and experimentally feasible alternative for generating large-amplitude SC states with high fidelity across a broad range of amplitudes $\alpha$, without requiring explicit photon addition or subtraction operations. However, we note that the success probability of our scheme is not high and is proportional to the square of the beam splitter deviation from the ideal $50:50$ configuration, i.e., $\epsilon{}^{2}$.
DISCUSSION
In this study, we have proposed a protocol for generating nG states by combining postselected WM with a weak cross-Kerr interaction. The core mechanism relies on injecting a single photon into an interferometer through the idler input port and heralding a successful event by its detection at a dark port. This process amplifies the single-photon nonlinearity via WVA, effectively transforming a given input signal state into a nG output without explicit photon-addition or -subtraction hardware.
Our theoretical analysis demonstrates that this versatile framework can generate a wide range of high-fidelity nG states. With coherent state inputs, the protocol produces states closely resembling SPAC states. For SV inputs, it yields high-fidelity TPASV states and their equivalents, squeezed number states. A particularly significant result is the protocol’s ability to act on nG inputs themselves: when applied to SC states, it not only enhances their non-Gaussianity, as evidenced by increased Wigner negativity, but also effectively amplifies small-amplitude cat states into large-amplitude ones ($\alpha\geq 2$) while maintaining fidelities exceeding 0.99, particularly in the case of odd SC states.
The primary advantage of our protocol is its versatility; a single apparatus can generate a continuous family of nG states, contrasting with methods that require specialized setups for each state type (2004; 2006). Its ability to amplify SC states addresses a key challenge in quantum optics, offering a distinct pathway compared to conventional photon operations that often degrade state quality at high amplitudes (PhysRevLett.101.233605; PhysRevA.110.023703).
This versatility comes with inherent practical challenges. The success probabilities for generating nG states lie between $10^{-6}$ and $10^{-3}$ for typical parameters ($\epsilon\thicksim 10^{-3}$, $g=0.01$), a direct trade-off intrinsically linked to the large WVA necessary for high-fidelity generation. For example, the heralded creation of an enlarged odd SC state with fidelity $F_{-}>0.96$ from an input with $\alpha=1.5$ occurs with a probability of $2\times 10^{-4}$. However, these probabilities fall within the practical regime enabled by high-repetition-rate laser sources (Ma_2015; XUE2024107048; CHEN2025111703), which can transform low per-trial probabilities into high absolute heralding rates of $10^{2}$ to $10^{4}$ states per second. This positions the predicted performance of our protocol on par with the demonstrated capabilities of other heralded nG state generation schemes, such as SPAC (2004) and cluster (PhysRevLett.97.110501) states, indicating that the achievable heralding rates would be on a similar scale.
Nevertheless, the protocol places stringent demands on component performance. Critically, the interferometer for the idler photon must exhibit high phase stability to maintain the precise destructive interference condition essential for the WVA effect; any instability introduces noise and degrades the fidelity of the output state. Furthermore, photon loss and non-ideal detector efficiency will introduce vacuum noise, reducing output purity. Additionally, the requirement of a weak nonlinearity ($g\langle n_{b}\rangle\ll 1$) constrains the usable interaction strength, making the achievement of a practical heralding rate a significant challenge.
Despite these limitations, our protocol establishes a valuable theoretical and practical framework. Its feasibility is supported by the widespread availability of high-repetition-rate laser sources, which are a standard tool in advanced quantum optics for generating the required input states and overcoming low probabilistic yields (PhysRevLett.101.250501; 2645; PhysRevA.110.023703). Looking forward, the convergence of this capability with continued progress in low-loss integrated photonics (PhysRevLett.133.083803), high-efficiency superconducting detectors (3803; Tao_2019; PhysRevApplied.19.064041), and engineered giant nonlinearities (2016) will directly enhance its performance and scalability, making it a promising pathway for advanced quantum state engineering.
In conclusion, we have shown that the synergistic use of WM and cross-Kerr nonlinearity provides a powerful method for nG state generation. Remarkably, this approach implicitly enacts nG operations, confirming the utility of WM as a tool for quantum state engineering. The ability to generate and amplify a continuous family of non-classical states from a unified framework opens up new possibilities for applications in CV quantum information processing.
METHODS
BS Transformation
The BS is a fundamental device in both classical and quantum optics, functioning to split a beam of light into transmitted and reflected components. The schematic of the BS used in our nG state generation protocol is shown in Fig. 11. It is well known that lossless two-mode BS (with two input and output ports) in quantum optics is described by the unitary matrix $U_{BS}$, which has the form
$\left(\begin{array}[]{c}a_{3}\\ a_{4}\end{array}\right)=U_{BS}(\theta,\varphi_{\tau},\phi_{\rho})\left(\begin{array}[]{c}a_{1}\\ a_{2}\end{array}\right),$ (29)
where the explicit expression of $U_{BS}$ is given in Eq. (2), and parameters $\theta,\varphi_{\tau}$ and $\phi_{\rho}$can be adjusted for specific purposes. Here, the annihilation operators for the first and second input modes are $a_{1}$ and $a_{2}$, respectively, and for the output ports are $a_{3}$ and $a_{4}$. The action of the BS transforms the states between the inner and outer paths of the interferometer. For example, if we consider a single-photon in mode $1$ and a vacuum state in mode $2$ of input ports, the BS transforming the state $|1\rangle_{1}|0\rangle_{2}$ to
$|\psi_{i}^{\prime}\rangle$ $=a_{1}^{{\dagger}}|0\rangle_{1}|0\rangle_{2}$
$=\left(e^{i\varphi_{\tau}}\cos\theta_{1}a_{3}^{{\dagger}}-e^{-i\phi_{\rho}}\sin\theta_{1}a_{4}^{{\dagger}}\right)|0\rangle_{3}|0\rangle_{4}$
$=e^{i\varphi_{\tau}}\cos\theta_{1}|1\rangle_{3}|0\rangle_{4}-e^{-i\phi_{\rho}}\sin\theta_{1}|0\rangle_{3}|1\rangle_{4}.$ (30)
The subindices of the states denote the corresponding paths in the interferometer. Consequently, by setting $\varphi_{\tau}=0$, and $\phi_{\rho}=\pi/2$,we obtain the preselection state $|\psi_{i}\rangle$ (see Eq. (3)) used in our protocol.
Furthermore, in the schematic of our protocol, a heralding event is defined by a click at detector $D_{1}$ and no click at $D_{2}$ . This corresponds to detecting a photon in one output port (e.g., arm 4) and no photon in the other (e.g., arm 3) of the second BS, which is characterized by the parameters $\theta_{2}$, $\varphi_{\tau}^{\prime}$ and $\phi_{\rho}^{\prime}$. This is equivalent to selecting the following path state inside the interferometer:
$|\psi_{f}^{\prime}\rangle$ $=a_{D1}^{{\dagger}}|0\rangle_{3}|0\rangle_{4}=a_{4}^{{\dagger}}|0\rangle_{3}|0\rangle_{4}$
$=(-e^{i\phi_{\rho}^{\prime}}\sin\theta_{2}a_{1}^{{\dagger}}+e^{i\varphi_{\tau}^{\prime}}cos\theta_{2}a_{2}^{{\dagger}})|0\rangle_{1}|0\rangle_{2}$
$=e^{i\varphi_{\tau}^{\prime}}cos\theta_{2}|0\rangle_{1}|1\rangle_{2}-e^{i\phi_{\rho}^{\prime}}\sin\theta_{2}|1\rangle_{1}|0\rangle_{2}.$ (31)
With the specific parameters $\theta_{2}$$\pi$/4, $\varphi_{\tau}^{\prime}=0,$ and $\phi_{\rho}^{\prime}=\pi/2$,the above state becomes the postselected state $|\psi_{f}\rangle$ given in Eq. (7).
Figure 11: Schematic representation of a BS. A BS scheme with two input ports and two output ports. The annihilation operators for the input modes are $a_{1}$ and $a_{2}$, and for the output modes are $a_{3}$ and $a_{4}$, respectively.
Numerical Simulation Details
All theoretical results and state characterizations were obtained through numerical simulations performed using the Quantum Toolbox in Python (QuTiP) (JOHANSSON20121760; JOHANSSON20131234). The simulation and analysis involved the following steps:
State Construction: The input states (coherent, squeezed vacuum, and Schrödinger cat states) were constructed using the relevant displacement ($D(\beta)$) and squeezing ( $S(r))$ operators within QuTiP.
Protocol Implementation: The final state $|\Phi_{f}\rangle$ was computed for each input state according to the derivation given in Results sections.
Fidelity Calculation: The fidelity between the generated state $\rho_{out}=|\Phi_{f}\rangle\langle\Phi_{f}|$ and a target pure state $|\psi_{target}\rangle$ was computed as $F=|\langle\psi_{target}|\Phi_{f}\rangle|^{2}$.
Wigner Function Computation: The Wigner function of a state $\rho$ was calculated using the standard formula $W(\alpha)=\frac{2}{\pi}Tr\left[D^{\dagger}(\alpha)\rho D(\alpha)(-1)^{b^{\dagger}b}\right].$ The Fock space dimension was set to 30 for simulations involving coherent, squeezed vacuum and cat states to ensure the computed states were well-contained within the Hilbert space and numerical truncation errors were negligible.
ACKNOWLEDGMENT
@@
AUTHOR CONTRIBUTIONS
Xiao-Xi Yao derived and simulated the theoretical results; Yusuf Turek wrote the main manuscript text and supervised the work . All authors participated in the discussion of results and writing and revision of the manuscript.
COMPETING INTERESTS
The authors declare no competing interests.
DATA AVAILABILITY
The source data that support the plots within this article and other findings of this study are publicly available at https://github.com/YaoXiaoxi023/code-and-data.
ADDITIONAL INFORMATION
Correspondence and requests for materials should be addressed to Yusuf Turek.